Three decision makers have assessed utilities for the following decision problem (payoff in dollars):

State of Nature
Decision Alternative S1 S2 S3
d1 30 40 -20
d2 80 100 -80

The indifference probabilities are as follows:

Indifference Probability (p)
Payoff Decision maker A Decision maker B Decision maker C
100 1.00 1.00 1.00
80 0.95 0.80 0.85
40 0.85 0.70 0.75
30 0.75 0.55 0.60
-20 0.60 0.25 0.50
-80 0.00 0.00 0.00

Find a recommended decision for each of the three decision makers, if P(s1) = 0.15, P(s2) = 0.70, and P(s3) = 0.15. (Note: For the same decision problem, different utilities can lead to different decisions.) If required, round your answers to two decimal places.

Decision maker A
EU(d1) =
EU(d2) =
Recommended decision:  
Decision maker B
EU(d1) =
EU(d2) =
Recommended decision:  
Decision maker C
EU(d1) =
EU(d2) =
Recommended decision:  
Answer #1

Decision maker A

Utility table

s1 s2 s3
d1 0.75 0.85 0.6
d2 0.95 1 0

EU(d1) = 0.15*0.75+0.70*0.85+0.15*0.6 = 0.7975

EU(d2) =  0.15*0.95+0.70*1+0.15*0 = 0.8425

Recommended decision : d2

Decision maker B

Utility table

s1 s2 s3
d1 0.55 0.7 0.25
d2 0.8 1 0

EU(d2) =  0.15*0.55+0.70*0.7+0.15*0.25 = 0.61

EU(d1) = 0.15*0.8+0.70*1+0.15*0= 0.82

Recommended decision : d2

Decision maker C

Utility table

s1 s2 s3
d1 0.6 0.75 0.5
d2 0.85 1 0

EU(d2) =  0.15*0.6+0.70*0.75+0.15*0.5= 0.69

EU(d1) = 0.15*0.85+0.70*1+0.15*0 = 0.8275

Recommended decision : d2